# LLM-based Test-driven Interactive Code Generation: User Study and Empirical Evaluation

### v1

@@ VIII-B Improving LLM code generation capabilities with verified test cases
Results of our benchmark evaluation indicate that, across all models studies, both implementations of TiCoder, TiCoder-PassFail and TiCoder-Output, can be used to augment the accuracy of an LLM through improved ranking and pruning. Specifically, we observe that using tests to better constrain the space of possible code suggestions can improve pass@k accuracy. This highlights the fact that current LLM capabilities may not be fully realized in practice: when prompted for multiple code suggestions, LLMs often are capable of generating a correct answer, but mechanisms to better rank the set of suggestions is needed. However, the performance boost provided by TiCoder is contingent on a set of high quality tests used by discriminative test ranking policy (Sec.  IV). If the underlying LLM is unable to generate high quality tests, the ranking and pruning mechanism may not be as helpful. Future work should explore more sophisticated mechanisms for generating high quality tests that capture important specifications about the code. In this work, we explore the user study scenario where TiCoder prunes away some, but not all of the AI generated code suggestions. It is worth noting that if the code suggestions generated by the underlying LLM predominantly exhibit consistent behavior, TiCoder can still be valuable to a user by providing meaningful tests alongside the code suggestions. For example, in a scenario where all of the suggestions are correct with respect to the user’s intent, TiCoder may not prune away any of the code suggestions, but provides some guarantees about program behaviour to the user.
@@ X Conclusion
In this work, we propose the workflow of test-driven interactive code generation using LLMs, and study it’s effectiveness through a user study and empirical evaluation on code generation benchmarks. Our findings provide encouraging results around guiding user intent clarification for generating more correct programs. Further algorithms and experiments exploring the TiCoder workflow can be found in [Lahiri et al.(2023)].
In future work, we plan to extend and evaluate our implementation reflecting real-world scenarios including: more complex programs, an in-situ user study for various software development tasks, and an empirical evaluation on realistic benchmarks such as CoderEval [Yu et al.(2023)] and NL2Fix [Fakhoury et al.(2023)].
Finally, we plan to explore if TiCoder can be extended to richer forms of formal specifications beyond tests, such as property based tests or pre- and post-conditions generated from user-defined prompts [Endres et al.(2023)].

### v2

@@ VIII-B Improving LLM code generation capabilities with verified test cases
Results of our benchmark evaluation indicate that, across all models studied, both implementations of TiCoder, TiCoder-PassFail and TiCoder-Output, can be used to augment the accuracy of an LLM through improved ranking and pruning. Specifically, we observe that using tests to better constrain the space of possible code suggestions can improve pass@k accuracy. This highlights the fact that current LLM capabilities may not be fully realized in practice: when prompted for multiple code suggestions, LLMs often are capable of generating a correct answer, but mechanisms to better rank the set of suggestions is needed. However, the performance boost provided by TiCoder is contingent on a set of high quality tests used by discriminative test ranking policy (Sec.  IV).
If the underlying LLM is unable to generate high quality tests, the ranking and pruning mechanism may not be as helpful. Future work should explore more sophisticated mechanisms for generating high quality tests that capture important specifications about the code. In this work, we explore the user study scenario where TiCoder prunes away some, but not all of the AI generated code suggestions.
It is worth noting that if the code suggestions generated by the underlying LLM predominantly exhibit consistent behavior, TiCoder can still be valuable to a user by providing meaningful tests alongside the code suggestions. For example, in a scenario where all of the suggestions are correct with respect to the user’s intent, TiCoder may not prune away any of the code suggestions, but provides some guarantees about program behaviour to the user.
VIII-C The Value of Execution Based Pruning
[…]
MBPP GPT-4-32k 81.56 78.26 87.81 80.88 91.58
HumanEval GPT-4-32k 76.10 65.25 81.37 70.39 86.99
TABLE V: Results of prompting GPT-4 with validation tests in the prompt
The TiCoder workflow helps users by automatically generating and ranking interesting test cases, instead of requiring the user to manually write them. However, users may have a set of test cases already in mind when querying a LLM. We explore how such a scenario compares to the TiCoder workflow by adding the set of validation tests, i.e. the ground truth test set provided in each dataset, to the code generation prompt. We instruct GPT-4-32k to generate a code suggestion that passes on the set of validation tests provided in the prompt. We then evaluate the $pass@k$ accuracy at $k=1$ and $k=5$ and compare to the TiCoder $pass@1@m$ accuracy from Table IV. Table V contains the result of this experiment, showing both experiments where a single test is added in the prompt, as well as an experiment where all available validation tests are added in the prompt, the exact number of tests varies across both datasets.
On both MBPP and HumanEval we observe a boost in $pass@k$ accuracy over the baseline prompt used Table IV. For example, on MBPP GPT-4-32k achieves a $pass@1$ of 67.13% with the baseline prompt (no tests) and 78.26% when including a single test in the prompt. This increases to a $pass@1$ of 80.88% when all tests are added in the prompt. As expected, this demonstrates that adding tests in the prompt does help improve model performance. However, recalling the $pass@1@1$ accuracy of the TiCoder-PassFail, within one user feedback loop, accuracy reaches 81.56%, out performing the $pass@1$ accuracy where a single test is included in the prompt by 3.31%. TiCoder even performs slightly better (0.68%) compared to the case when all tests are in the prompt.
This indicates that even if the user supplies all test, there is no guarantee that the underlying LLM will fulfill the user intent and generate a code suggestion that passes on the tests. Conversely, with only one user interaction on a highly distinguishing test, obtained by ranking LLM generated tests, code generation accuracy greatly improves, matching the scenario where several tests are provided to the LLM without the added burden on the user to manually come up with test cases.
VIII-D Considerations for AI-Generated Tests: Precondition Violations
Tests generated by LLMs may contain pre-condition violating inputs that would cause the reference implementation to crash or fail, and would result in a Undefined response from the user, resulting in no pruning of code.
As an illustrative case, consider from MBPP, the example of a reference implementation with (implicit) precondition that the argument nums array is non-empty, and throws a divisionByZero at the return statement for an empty nums array.
1from array import array
2def zero_count(nums):
3 """Write a function to find the ratio of zeroes to non-zeroes in an array of integers."""
4 n = len(nums)
5 n1 = 0
6 for x in nums:
7 if x == 0:
8 n1 += 1
9 else:
10 None
11 return n1/(n-n1)
The following test is ranked highest by the discriminative ranking strategy:
1assert zero_count([]) == 0, "Empty List"
For this test, out of the (deduplicated) 80 code suggestions from code-davinci-002, 8 suggestions pass the test, 11 suggestions fail the test and 61 suggestions crash on this test. The score for this test is 8/11 = 0.73, which is the highest among all tests. However, this test does not lead to any code pruning as the user responds Undefined in both the TiCoder-PassFail and TiCoder-Output scenarios since the empty array causes the reference implementation above to fail. Thus, our results in RQ2 account for tests with Undefined responses, reflecting the possible real world impact of pre-condition violating tests on the accuracy boost provided by TiCoder.
@@ X Conclusion
In this work, we propose the workflow of test-driven interactive code generation using LLMs, and study it’s effectiveness through a user study and empirical evaluation on code generation benchmarks. Our findings provide encouraging results around guiding user intent clarification for generating more correct programs.
In future work, we plan to extend and evaluate our implementation reflecting real-world scenarios including: more complex programs, an in-situ user study for various software development tasks, and an empirical evaluation on realistic benchmarks such as CoderEval [47] and NL2Fix [43]. Finally, we plan to explore if TiCoder can be extended to richer forms of formal specifications beyond tests, such as property based tests or pre- and post-conditions generated from user-defined prompts [48].